# Spatially-Augmented Sequence-to-Sequence Neural Diarization for Meetings

### v1

@@ Abstract
Index Terms—  Speaker Diarization, Online Speaker Diarization, Sequence-to-Sequence Neural Diarization
@@ 1 Introduction
Speaker diarization aims to answer the “who-spoke-when” question [1], serving as a fundamental pre-processing step for downstream tasks (e.g., speech recognition) [2]. Despite notable progress, speaker diarization in meetings remains challenging due to overlapping speech, unreliable speaker embeddings, reverberation, etc.
@@
Early studies on speaker diarization focused on modularized approaches [3, 4, 5], which segment audio into short utterances and cluster them by speaker embedding similarity. Assuming each segment contains only one speaker, these methods perform poorly on overlaps. End-to-End Neural Diarization (EEND) addresses this by casting diarization as a multi-label prediction task, achieving greater robustness [6, 7, 8]. More recently, Target-Speaker Voice Activity Detection (TSVAD) combines modularized and neural methods with strong performance [9, 10, 11, 12, 13, 14]. Sequence-to-sequence diarization (S2SND) further advances online diarization [15]. However, most approaches still rely solely on acoustic embeddings, which are often unreliable in real meetings. This raises a key question: how can spatial cues from multi-channel recordings improve diarization?
@@
In multi-channel processing, spatial cues can be integrated in three ways: 1) speech enhancement such as beamforming to generate cleaner inputs [16, 17], though often introducing distortions that harm speaker discrimination; 2) channel-fusion or attention modules that aggregate signals [18], but typically perform blind fusion rather than true localization; and 3) explicit features like direction-of-arrival (DOA) estimates [19]. Among them, explicit DOA provides direct directional evidence to separate simultaneous speakers and thus holds greater promise for meeting diarization.
@@
To effectively integrate DOA cues into diarization, two challenges must be solved: robust extraction of high-quality DOA under noise, reverberation, and multi-source conditions, and effective fusion of these cues into diarization models for both online and offline use. Traditional localization methods (e.g., statistical filtering [19], HMM-based clustering [20], steered response power (SRP) [21]) have been studied, but recent deep learning approaches show greater potential [22, 23, 24]. In particular, SRP-DNN [24] learns direct-path phase differences and builds an enhanced SRP spectrum with iterative peak detection/removal, achieving robust DOA estimation in adverse conditions. We therefore adopt SRP-DNN for DOA extraction and S2SND [15] as the diarization backbone.
@@
Building on these observations, we propose SA-S2SND (Spatially-Augmented S2SND). It injects DOA cues from SRP-DNN as explicit auxiliary inputs into the S2SND backbone, enhancing discriminability and improving performance in both online and offline modes. To improve generalization and decouple spatial cues from specific arrays, we introduce a simulated DOA strategy: real multi-channel speech is paired with estimated DOA, while simulated multi-channel speech (from single-channel data) is paired with simulated DOA. This reduces reliance on large multi-channel corpora and improves adaptability to diverse arrays and conditions. Our main contributions are as follows:
@@ 2 Methods
In this section, we introduce the methodology with two components: (1) revisiting direction-of-arrival (DOA) estimation from multi-channel inputs using SRP-DNN [24]; and (2) integrating these DOA cues into the sequence-to-sequence neural diarization (S2SND) backbone [15] to form our proposed SA-S2SND framework.
@@ 2.1 DOA Estimation with SRP-DNN
The first component is the direction-of-arrival (DOA) estimation module. We adopt SRP-DNN [24], which estimates robust DOAs for multiple sources by learning direct-path phase differences (DP-IPDs) with a causal CRNN. For the $k$-th source, DOA is represented as $\boldsymbol{\theta}_{k}=\left[\theta_{k}^{\mathrm{ele}},\theta_{k}^{\mathrm{azi}}\right]^{T}$, with elevation $\theta_{k}^{\mathrm{ele}}\in[0,\pi]$ and azimuth $\theta_{k}^{\mathrm{azi}}\in[-\pi,\pi)$.
Given $M$ microphones, the network extracts multi-channel short-time features (log-magnitude and phase of microphone pairs) and outputs a summed DP-IPD vector $\widehat{R}_{mm^{\prime}}(n)$ for each frame $n$. The training target is a weighted sum of direct-path IPD vectors:
$R_{mm^{\prime}}(n)=\sum_{k=1}^{K}\beta_{k}(n)\,r_{mm^{\prime}}(\boldsymbol{\theta}_{k}(n)),$ (1)
where $r_{mm^{\prime}}(\boldsymbol{\theta})$ is the DP-IPD vector of candidate DOA and $\beta_{k}(n)\in[0,1]$ is the activity probability of source $k$. This avoids permutation issues, and the network minimizes the MSE between $\widehat{R}_{mm^{\prime}}(n)$ (prediction) and $R_{mm^{\prime}}(n)$. From $\{\widehat{R}_{mm^{\prime}}(n)\}$, an SRP-style spatial spectrum is constructed:
$P^{\prime}(\boldsymbol{\theta};n)=\frac{2}{M(M-1)F}\sum_{m=1}^{M-1}\sum_{m^{\prime}=m+1}^{M}\Re\!\{\widehat{R}_{mm^{\prime}}(n)^{\mathsf{H}}r_{mm^{\prime}}(\boldsymbol{\theta})\},$ (2)
@@
To localize multiple sources, SRP-DNN applies an iterative detection-and-removal (IDL) strategy for each frame $n$: 1) obtain $\widehat{R}_{mm^{\prime}}(n)$ via the CRNN; 2) construct $P^{\prime}(\boldsymbol{\theta};n)$ and find the dominant peak as candidate DOA; 3) estimate its weight, namely the energy ratio of $\widehat{R}_{mm^{\prime}}(n)$; 4) remove the source’s contribution from $\widehat{R}_{mm^{\prime}}(n)$ and repeat until the predefined threshold is reached. This yields cleaner spatial spectra and more reliable DOA estimates under reverberant, noisy, and multi-speaker conditions, providing explicit spatial cues for the diarization backbone.
@@ 2.2 Proposed SA-S2SND
In this work, we propose SA-S2SND, extending our sequence-to-sequence diarization (S2SND) [15] and its multi-channel variant (MC-S2SND) [25] with cross-channel attention. Although these models support both online and offline diarization, they lack explicit spatial information essential for separating simultaneous speakers in meetings. SA-S2SND addresses this by injecting DOA cues from SRP-DNN into the backbone.
@@ 2.2.1 Architecture
The SA-S2SND architecture is shown in Fig. 1. Its backbone follows S2SND [15], which formulates diarization as a sequence-to-sequence task of joint speaker detection and speaker representation, consisting of four modules: an extractor, an encoder, and two coupled decoders.
The extractor is a ResNet [26] with segmental statistical pooling (SSP) [11], producing frame-level embeddings. A Conformer encoder [27] models long-range dependencies and outputs contextualized features $\hat{X}$. On top, two symmetric decoders operate: 1) the representation decoder (Rep.) uses extractor outputs and voice-activity queries to generate target embeddings $\widehat{E}$; 2) the detection decoder (Det.) uses encoder features $\hat{X}$ and embedding queries to predict activities $\widehat{Y}$. These decoders form an inverse pair, conditioning embeddings on activities and vice versa, thus avoiding clustering or PIT assignment.
Building on this backbone, SA-S2SND incorporates spatial cues from SRP-DNN. For each block, the DOA model outputs azimuths of active speakers, $\theta_{k}^{\mathrm{azi}}\in\mathbb{R}^{{T^{\prime\prime}}\times\hat{N}}$ with $\hat{N}\leq 2$. Since SRP-DNN produces azimuths in $180^{\circ},180^{\circ})$ at $5^{\circ}$ resolution, we construct a DOA matrix $\mathbf{O}\in\mathbb{R}^{{T^{\prime\prime}}\times A}$, where each row encodes the probability of activity at a given azimuth, aligned with $\beta$ in Eq. (1). Because SRP-DNN runs at coarser resolution ($T^{\prime\prime}$) than frame-level embeddings ($T$), $\mathbf{O}$ is upsampled by nearest-neighbor interpolation, projected to dimension $D$, and fused with acoustic embeddings via residual addition:
@@
where $D$ is the hidden dimension. This enriches encoder representations with explicit spatial cues, analogous to positional encoding, improving speaker discriminability.
In the multi-channel case, a cross-channel attention module is applied to extractor outputs, following MC-S2SND [25]. While MC-S2SND mainly aggregates frame-level speaker embeddings from different channels without preserving explicit spatial information. In contrast, SA-S2SND injects DOA cues as directional information, providing complementary spatial evidence and further improving performance.
@@ 2.2.2 Training Process
Part A: Single-channel model + Multi-channel DOA (Stages 1–3). The model takes single-channel audio, while DOA features are provided in parallel: SRP-DNN estimates DOA for real recordings, and pseudo-DOA is generated online for simulated mixtures using VAD with random azimuth assignment and perturbation.
@@
  • Stage 3. Fine-tune the entire single-channel SA-S2SND with reduced LR=1e-5.
Part B: Multi-channel model + Multi-channel DOA (Stages 4–5). The model is upgraded with cross-channel attention (as in MC-S2SND) to process multi-channel audio, while DOA is still provided by SRP-DNN.
@@ 2.2.3 Inference Process
The inference of SA-S2SND follows the block-wise sliding-window scheme of S2SND [15]. Each block has left, chunk, and right contexts with lengths $L_{\text{left}}$, $L_{\text{chunk}}$, and $L_{\text{right}}$, giving $L=L_{\text{left}}+L_{\text{chunk}}+L_{\text{right}}$. Setting chunk shift to $L_{\text{chunk}}$ yields latency $L_{\text{chunk}}+L_{\text{right}}$.
@@
After the first (online) pass, the final embedding buffer enables re-decoding of the full recording, where acoustic and DOA-enhanced representations jointly improve offline results. The inference follows S2SND [15], with DOA injection as the only modification.
Table 1: Performance on Alimeeting test sets with various model size and inferring conditions. The Diarization Error Rates (DERs) are reported without Oracle VAD and collar tolerance. Ch. Num. means the number of channels for S2SND
[b] ID Model Size Ch. Num. Usage of DOA Methods 1-2 SPKs 2+ SPKs Total Online DER (%) Offline DER (%) Online DER (%) Offline DER (%) Online DER (%) Offline DER (%) E1 Small(16.56M) 1 S2SND 6.41 5.41 20.72 17.58 16.03 13.59 E2 Small(16.56M+0.86M) 1 SA-S2SND 6.35 5.40 19.75 16.10 15.35 12.59 E3 Small(18M) 8 S2SND 5.86 5.29 19.24 16.44 14.85 12.79 E4 Small(18M+0.86M) 8 SA-S2SND 5.90 5.01 16.33 13.68 12.93 10.84 E5 Medium(45.96M) 1 S2SND 6.23 5.46 18.57 16.06 14.53 12.59 E6 $\approx$ 69M 1 BUT System [28] $\dagger$} - - - - - 12.50
  • The lowest online and offline DERs (in our experiments) are highlighted by the gray background.
  • $\dagger$}
    State-of-the-art by June 2025.
@@ 3.1 Datasets
For simulated data, we use VoxCeleb2 [29] (1M utterances, 6,112 speakers) with on-the-fly mixture generation. Real data come from AliMeeting [30] (104.75h train, 4h eval, 10h test) containing 8-channel far-field array and headset recordings. We use the far-field array signals after dereverberation with NARA-WPE  [^1] https://github.com/fgnt/nara_wpe. The model is trained on the training set, validated on the evaluation set, and tested without oracle VAD or collar tolerance.
@@ 3.2 Network Configurations
The network has five components: extractor, channel-attention, encoder, DOA model, and two decoders. We implement S2SND-Small and S2SND-Medium, both using a ResNet-34 [26] extractor. The small version uses residual widths {32,64,128,256}, and the medium {64,128,256,512}. The channel-attention module is a 2-block Transformer [31] with 512-dim 8-head attention and 1024-dim feed-forward layers.
The encoder is a Conformer [27] with kernel size 15, and both decoders follow the Speaker-wise Decoder (SW-D) of S2SND [15]. In S2SND-Small, we use 256-dim 8-head attention with 512-dim feed-forward layers, and S2SND-Medium uses 384-dim with 768-dim feed-forward layers. All have 4 blocks, with other settings identical. In total, S2SND-Small and S2SND-Medium have 16.56M and 45.96M parameters, respectively. The single- and multi-channel versions differ only by the added channel-attention module. The DOA model (0.86M) is used only for estimation and not updated during training.
@@ 3.3 Training and Inference Details
For training, the audio is segmented into 8s windows with 2s overlaps, which are then normalized and converted into 80-dimensional log-Mel filterbanks (25-ms frame length, 10-ms shift). The output resolution is 10 ms [15], with a maximum of $N=30$ speakers. To ensure permutation invariance, input embeddings and activities are randomly shuffled with reassigned labels. Data augmentation includes Musan noise [32] and RIR reverberation [33]. Optimization uses AdamW [34] with BCE and ArcFace losses [35] ($s=32,m=0.2$). Experiments run on two RTX-A6000 GPUs.

### v2

# Li Cheng Liu Li

@@ Abstract
keywords:
Speaker Diarization, Online Speaker Diarization, Sequence-to-Sequence Neural Diarization
@@ 1 Introduction
Speaker diarization aims to answer the ``who-spoke-when'' question [park2022review], serving as a fundamental pre-processing step for downstream tasks (e.g., speech recognition) [kanda2020joint]. Despite notable progress, speaker diarization in meetings remains challenging due to overlapping speech, unreliable speaker embeddings, reverberation, etc.
@@
Early studies on speaker diarization focus on modularized pipelines [wang2018speaker, landini2022bayesian], which segment audio and cluster them by speaker embedding similarity. Assuming each segment contains only one speaker, these methods perform poorly on overlaps. End-to-End Neural Diarization (EEND) addresses this by formulating diarization as a multi-label prediction task, achieving greater robustness [fujita2019end_1, horiguchi2022encoder]. More recently, Target-Speaker Voice Activity Detection (TSVAD) combines modular and neural methods with strong performance [medennikov2020target, cheng2023target, wang2022similarity, wang2021dku], while Sequence-to-Sequence Neural Diarization (S2SND) further advances online diarization [cheng2024sequence]. However, most approaches still rely solely on acoustic embeddings, which are often unreliable in real meetings. In contrast, spatial cues provide an orthogonal source of information, since speakers typically occupy different physical locations. This raises a key question: how can spatial cues from multi-channel recordings improve diarization?
@@
In multi-channel processing, spatial cues can be integrated in three ways: 1) speech enhancement such as beamforming to generate cleaner inputs [Anguera, Boeddeker], though they may often introduce distortions that harm speaker discrimination; 2) channel-fusion or attention modules that aggregate signals [Horiguchi2022], but typically perform blind fusion rather than true localization; and 3) explicit features like direction-of-arrival (DOA) estimates [doa2008]. Among them, explicit DOA provides direct directional evidence to separate simultaneous speakers and thus holds greater promise for meeting diarization.
@@
To effectively integrate DOA cues into diarization, two challenges must be solved: robust extraction of high-quality DOA under noise, reverberation, and multi-source conditions, and effective fusion of these cues into diarization models for both online and offline scenarios. Traditional localization methods (e.g., statistical filtering [doa2008], HMM-based clustering [doa2021], steered response power (SRP) [srp]) have been studied, but recent deep learning approaches show greater potential [nguyen2020robust, 9413923, 9746624]. In particular, SRP-DNN [9746624] learns direct-path phase differences and builds an enhanced SRP spectrum with iterative peak detection, achieving robust multi-speaker DOA estimation in adverse conditions. We therefore adopt SRP-DNN for DOA extraction and S2SND [cheng2024sequence] as the diarization backbone.
@@
Building on these observations, we propose SA-S2SND (Spatially-Augmented S2SND). It injects DOA cues from SRP-DNN as explicit auxiliary inputs into the S2SND backbone, enhancing discriminability and improving performance in both online and offline modes. To improve generalization and decouple spatial cues from specific arrays, we introduce a simulated DOA strategy: real multi-channel speech is paired with estimated DOA, while simulated multi-channel speech (from single-channel data) is paired with simulated DOA. This reduces reliance on large-scale multi-channel corpora and improves adaptability to diverse arrays and conditions. Our main contributions are as follows:
@@ 2 Methods
In this section, we introduce the method with two components: (1) revisiting direction-of-arrival (DOA) estimation from multi-channel inputs using SRP-DNN [9746624]; and (2) integrating these DOA cues into the sequence-to-sequence neural diarization (S2SND) backbone [cheng2024sequence] to form our proposed SA-S2SND framework.
@@ 2.1 DOA Estimation with SRP-DNN
The first component is the direction-of-arrival (DOA) estimation module. We adopt SRP-DNN [9746624], which estimates robust DOAs for multiple sources by learning direct-path inter-channel phase differences (DP-IPDs) with a causal CRNN. This model is lightweight (0.86M) and easy to retrain for different array configurations. For the $k$-th source, DOA is represented as $\bm{\theta}_{k}=\left[\theta_{k}^{\mathrm{ele}},\theta_{k}^{\mathrm{azi}}\right]^{T}$, with elevation $\theta_{k}^{\mathrm{ele}}\in[0,\pi]$ and azimuth $\theta_{k}^{\mathrm{azi}}\in[-\pi,\pi)$.
Specifically, given $M$ microphones, the network extracts multi-channel short-time features (log-magnitude and phase of microphone pairs) and outputs a summed DP-IPD vector $\widehat{R}_{mm^{\prime}}(n)$ for each frame $n$. The training target is a weighted sum of direct-path IPD vectors:
$R_{mm^{\prime}}(n)=\sum_{k=1}^{K}\beta_{k}(n)\,r_{mm^{\prime}}(\bm{\theta}_{k}(n)),$ (1)
where $r_{mm^{\prime}}(\bm{\theta})$ is the DP-IPD vector of candidate DOA and $\beta_{k}(n)\in[0,1]$ is the activity probability of source $k$. This avoids permutation issues, and the network minimizes the MSE between $\widehat{R}_{mm^{\prime}}(n)$ (prediction) and $R_{mm^{\prime}}(n)$. From $\{\widehat{R}_{mm^{\prime}}(n)\}$, an SRP-style spatial spectrum is constructed:
$P^{\prime}(\bm{\theta};n)=\frac{2}{M(M-1)F}\sum_{m=1}^{M-1}\sum_{m^{\prime}=m+1}^{M}\Re\!\{\widehat{R}_{mm^{\prime}}(n)^{\mathsf{H}}r_{mm^{\prime}}(\bm{\theta})\},$ (2)
@@
To localize multiple sources, SRP-DNN applies an iterative detection-and-removal (IDL) strategy for each frame $n$: 1) obtain $\widehat{R}_{mm^{\prime}}(n)$ via the CRNN; 2) construct $P^{\prime}(\bm{\theta};n)$ and find the dominant peak as candidate DOA; 3) estimate its weight, namely the energy ratio of $\widehat{R}_{mm^{\prime}}(n)$; 4) remove the source’s contribution from $\widehat{R}_{mm^{\prime}}(n)$ and repeat until the predefined threshold is reached. This yields cleaner spatial spectra and more reliable DOA estimates under reverberant, noisy, and multi-speaker conditions, providing explicit spatial cues for the diarization backbone.
@@ 2.2 Proposed SA-S2SND
In this work, we propose SA-S2SND, extending our sequence-to-sequence diarization (S2SND) [cheng2024sequence] and its multi-channel variant (MC-S2SND) [cheng25b_interspeech] with cross-channel attention. Although these models support both online and offline diarization, they lack explicit spatial information that essential for separating simultaneous speakers in meetings. SA-S2SND addresses this by injecting DOA cues from SRP-DNN into the backbone.
@@ 2.2.1 Architecture
The SA-S2SND architecture is shown in Fig. 1. Its backbone follows S2SND [cheng2024sequence], which formulates diarization as a sequence-to-sequence task of joint speaker detection and speaker representation, consisting of four modules: an extractor, an encoder, and two coupled decoders.
The extractor is a ResNet [he2016deep] with segmental statistical pooling (SSP) [wang2022similarity], producing frame-level embeddings. A Conformer encoder [gulati2020conformer] models long-range dependencies and outputs contextualized features $\hat{X}$. On top, two symmetric decoders operate: 1) the representation decoder (Rep.) uses extractor outputs and voice-activity queries to generate target embeddings $\widehat{E}$; 2) the detection decoder (Det.) uses encoder features $\hat{X}$ and embedding queries to predict activities $\widehat{Y}$. These decoders form an inverse pair, conditioning embeddings on activities and vice versa, thus avoiding clustering or PIT assignment.
Building on this backbone, SA-S2SND incorporates spatial cues. For each block, the DOA model outputs azimuths of active speakers, $\theta_{k}^{\mathrm{azi}}\in\mathbb{R}^{{T^{\prime\prime}}\times\hat{N}}$ with $\hat{N}\leq 2$. Since SRP-DNN produces azimuths in $180^{\circ},180^{\circ})$ at $5^{\circ}$ resolution, we construct a DOA matrix $\mathbf{O}\in\mathbb{R}^{{T^{\prime\prime}}\times A}$, where each row encodes the probability of activity at a given azimuth, aligned with $\beta$ in Eq. (1). Because SRP-DNN runs at coarser resolution ($T^{\prime\prime}$) than frame-level embeddings ($T$), $\mathbf{O}$ is upsampled by nearest-neighbor interpolation, projected to dimension $D$, and fused with acoustic embeddings via residual addition:
@@
where $D$ is the hidden dimension. This enriches encoder representations with directional priors, analogous to positional encoding, enabling the model to distinguish speakers that overlap temporally but originate from different spatial locations.
In the multi-channel case, a cross-channel attention module is applied to extractor outputs, following MC-S2SND [cheng25b_interspeech]. While MC-S2SND aggregates frame-level speaker embeddings from different channels without preserving explicit spatial information. In contrast, SA-S2SND injects DOA cues as directional information, providing complementary spatial evidence and further improving performance.
@@ 2.2.2 Training Process
Part A: Single-channel model + Multi-channel DOA (Stages 1–3). The model takes single-channel audio, while DOA features are provided in parallel: SRP-DNN estimates DOA for real recordings, and pseudo-DOA is generated online for simulated mixtures using VAD with random DOA assignment and perturbation.
@@
  • Stage 3. Fine-tune the entire single-channel SA-S2SND with a reduced LR of 1e-5.
Part B: Multi-channel model + Multi-channel DOA (Stages 4–5). The model is upgraded with cross-channel attention to process multi-channel audio, while DOA is still provided by SRP-DNN.
@@ 2.2.3 Inference Process
The inference of SA-S2SND follows the block-wise sliding-window scheme of S2SND [cheng2024sequence]. Each block has left, chunk, and right contexts with lengths $L_{\text{left}}$, $L_{\text{chunk}}$, and $L_{\text{right}}$, giving $L=L_{\text{left}}+L_{\text{chunk}}+L_{\text{right}}$. Setting chunk shift to $L_{\text{chunk}}$ yields latency $L_{\text{chunk}}+L_{\text{right}}$.
@@
After the first (online) pass, the final embedding buffer enables a second-pass offline decoding, where acoustic and DOA-enhanced representations jointly improve offline results. The inference follows S2SND [cheng2024sequence], with DOA injection as the only modification.
@@ 3.1 Datasets
For simulated data, we use VoxCeleb2 [Chung18b] (1M utterances, 6,112 speakers) with on-the-fly mixture generation. Real data come from AliMeeting [alimeeting] (104.75h train, 4h eval, 10h test) containing 8-channel far-field array and headset recordings. We use the far-field array signals after dereverberation with NARA-WPE  [^1] https://github.com/fgnt/nara_wpe. Additionally, we also use a compound dataset for real data: Alimeeting [alimeeting], DIHARD III [ryant2020third], VoxConverse [voxconverse], MISP2022 [chen2022misp]. The model is trained on the training set, validated on the evaluation set, and tested without oracle VAD or collar tolerance.
@@
Table 1: Overlap ratio of more than two speakers across datasets.
   Dataset    Ratio (%)
   AliMeeting    6.0
   AISHELL-4    0.5
   AMI    3.0
Table 2: Performance on Alimeeting test sets with various model size and inferring conditions. The Diarization Error Rates (DERs) are reported without Oracle VAD and collar tolerance. Ch. Num. means the number of channels for S2SND.
[…]
    We use the compound dataset. The lowest online and offline DERs are highlighted by the gray background.
@@ 3.2 Network Configurations
The network has five components: extractor, channel-attention, encoder, DOA model, and two decoders. We implement S2SND-Small and S2SND-Medium, both using a ResNet-34 [he2016deep] extractor. The small version uses residual widths {32,64,128,256}, and the medium {64,128,256,512}. The channel-attention module is a 2-block Transformer [vaswani2017attention] with 512-dim 8-head attention and 1024-dim feed-forward layers. The encoder is a Conformer [gulati2020conformer] with a kernel size of 15, and both decoders follow the speaker-wise decoder of S2SND [cheng2024sequence]. And the two model variants share the same 4-block structure but differ in their hidden dimensions: S2SND-Small (256-dim, 8-head attention; 512-dim FFN) and S2SND-Medium (384-dim attention; 768-dim FFN), totaling 16.56M and 45.96M parameters, respectively. The single- and multi-channel versions differ only by the added channel-attention module. The DOA model (0.86M) is used only for estimation and not updated during training.
@@ 3.3 Training and Inference Details
For training, the audio is segmented into 8s windows with 2s overlaps, which are then normalized and converted into 80-dimensional log-Mel filterbanks (25-ms frame length, 10-ms shift). The output resolution is 10 ms [cheng2024sequence], with a maximum of $N=30$ speakers. To ensure permutation invariance, input embeddings and activities are randomly shuffled with reassigned labels. Data augmentation includes Musan noise [snyder2015musan] and RIR reverberation [ko2017study]. Optimization uses AdamW [loshchilov2017decoupled] with BCE and ArcFace losses [deng2019arcface] ($s=32,m=0.2$). Experiments run on two RTX-A6000 GPUs.